OpenStudy (anonymous):

l'Hospital's rule: sin(x)/3x

OpenStudy (anonymous):

i need to find limits!

zepdrix (zepdrix):

as x approaches 0?

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

You don't need L'Hop for this one. Did they instruct you to use it though? :)

OpenStudy (anonymous):

well they said if we didn't have then we don't but we are encouraged to? Can you show me both ways?

OpenStudy (anonymous):

no need

OpenStudy (anonymous):

so do you get the derivatives of both? if I was to use LHop:?

zepdrix (zepdrix):

The first way is to remember a really important identity, it shows up a lot to try to commit it to memory.\[\large \lim_{x \to 0}\frac{\sin x}{x}=1\]

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{1}{3}\frac{\sin(x)}{x}\]

OpenStudy (anonymous):

\[=\frac{1}{3}\lim_{x\to 0}\frac{\sin(x)}{x}\]

OpenStudy (anonymous):

ok~

OpenStudy (anonymous):

what @zepdrix said finishes it

OpenStudy (anonymous):

so if I was to use L'hop rule????

zepdrix (zepdrix):

Ok let's see :)

OpenStudy (anonymous):

or what happens to the 1/3? when the method you showed me is used?

OpenStudy (anonymous):

it just sits there

OpenStudy (anonymous):

lol so the final answer is?

OpenStudy (anonymous):

\[\lim_{x\to a}cf(x)=c\lim_{x\to a}f(x)\]

zepdrix (zepdrix):

\[\large \color{orangered}{\lim_{x \to 0}\frac{\sin x}{x}=1}\] \[\large \frac{1}{3}\color{orangered}{\lim_{x\to 0}\frac{\sin(x)}{x}} \qquad=\qquad \frac{1}{3}\color{orangered}{1}\]

OpenStudy (anonymous):

ahh...

OpenStudy (anonymous):

oooh orange is the new black!

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

\[\large \lim_{x \to 0}\frac{\sin x}{3x}\qquad\rightarrow\qquad \frac{0}{0}\] This limit is approaching 0/0, which IS one of the indeterminate forms we're allowed to apply L'Hopital's Rule to! So taking the derivative of the top and bottom (separately) gives us,\[\large L'H \qquad=\qquad \lim_{x \to 0}\frac{\cos x}{3}\]

OpenStudy (anonymous):

right.

zepdrix (zepdrix):

So what happens as x approaches 0 now? :O hmm

OpenStudy (anonymous):

and cos(x) equals one so 1/3?

zepdrix (zepdrix):

yay good job \:D/

zepdrix (zepdrix):

cos(0)=1, yes

OpenStudy (anonymous):

ok can we work on one more?

OpenStudy (anonymous):

so I can make sure I got this?

zepdrix (zepdrix):

sure :3

OpenStudy (anonymous):

cos(x)/ sinx+1

OpenStudy (anonymous):

do we need to use LHop?

zepdrix (zepdrix):

x approaching?

OpenStudy (anonymous):

oh zero lol

zepdrix (zepdrix):

Hmmm.. Can't we just evaluate this one directly? Plugging in x=0 gives us,\[\large \frac{\cos0}{\sin0+1} \qquad=\qquad \frac{1}{0+1}\]

OpenStudy (anonymous):

i don't know I guess...that'll be ok? hopefully?

zepdrix (zepdrix):

Well I worked it out with L'Hopital and it's giving us the same answer. So yah.. it should be ok. Not sure why you were given an easy problem like that. Maybe just to confuse you lol.

OpenStudy (anonymous):

hmmm ok! ONE MORE! just one more?

zepdrix (zepdrix):

fine fine fine XD

OpenStudy (anonymous):

xe^-x

OpenStudy (anonymous):

as x approaches infinity

OpenStudy (anonymous):

hmph i hate e's

zepdrix (zepdrix):

Ok so you'll want to make sure you get comfortable with exponents. This can be a little tricky to evaluate directly if you don't remember how negative exponents work. Let's look at the exponential a sec, \(\large \lim_{x\to\infty}e^{-x}\qquad\to\qquad0\) Understand why this is approaching 0?

OpenStudy (anonymous):

no. lol

OpenStudy (anonymous):

i thought it was approaching infinity?

OpenStudy (anonymous):

oh shoot I gotta run! but can you please explain it to mee and I'll def check it when I get back??

zepdrix (zepdrix):

Let's plug in a really really big value (close to infinity) for x and see what happens. \[\huge e^{-99999} \qquad=\qquad \frac{1}{e^{99999}} \qquad\approx\qquad 0\]Remember how the negative on the exponent tells us to flip it to the denominator? So this gives us 1/(a bajillion), which is really really close to zero.

zepdrix (zepdrix):

fine fine fine, ill leave some notes -_-

OpenStudy (anonymous):

I REALLY like how your explaining it!!!

OpenStudy (anonymous):

hahaa thanks man

OpenStudy (anonymous):

ohhh and i got ur last note. I understand why its approaching zero

zepdrix (zepdrix):

\[\large \lim_{x\to\infty}xe^{-x} \qquad\to\qquad \infty\cdot0\] This is an indeterminate form, but it's NOT in the proper form for us to apply L'Hopital's Rule. So like I did in that example with the negative exponent, let's do it here also.\[\large \lim_{x\to\infty}xe^{-x} \qquad=\qquad \lim_{x\to\infty}\frac{x}{e^x} \qquad\to\qquad \frac{\infty}{\infty}\] Ok good! Now it's giving us the form we need for L'Hopital.

zepdrix (zepdrix):

Applying L'Hopital's Rule, separately taking the derivative of the top and bottom, gives us,\[\large L'Hop \qquad=\qquad \lim_{x\to\infty}\frac{1}{e^x}\]

zepdrix (zepdrix):

If we want to, we could rewrite this as,\[\large \lim_{x\to\infty}\frac{1}{e^x}\qquad=\qquad\lim_{x\to\infty}e^{-x}\] And I think we already talked about this particular limit before :) So just scroll up if you forgot. ( The thing with the 1/[a bajillion] ).

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

haha i just saw the reply from the other person!